OpenStudy (anonymous):

Let u = <-6, 1>, v = <-5, 2>. Find -4u + 2v (a)<44, -12> (b)<34, -8> (c)<14, 0> (d)<14, 3>

OpenStudy (anonymous):

@Butterfield1215 help please?

OpenStudy (butterfield1215):

i have no idea what this im so sorry

TheSmartOne (thesmartone):

\(\sf k<u,v>~ = ~<ku,kv>\)

OpenStudy (anonymous):

What would K represent? @TheSmartOne

OpenStudy (phi):

-4u k is the -4

TheSmartOne (thesmartone):

\(\sf\large <v_1, v_2> + <u_1,u_2> ~= ~<v_1+u_1, v_2+u_2>\)

OpenStudy (anonymous):

11+3=14?

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (phi):

where did you get 11+3 ? -4 times each number in u +2 times each number in v then add corresponding components to get a new vector

OpenStudy (anonymous):

so -4(-6)=-24 -4(1)=-4 (24,-4) 2(-5)=-10 2(2)=4 (-10,4)

OpenStudy (phi):

yes, so far so good

TheSmartOne (thesmartone):

negative number times a negative number = a positive number

TheSmartOne (thesmartone):

So what is \(\sf -4 \times -6 = ?\)

TheSmartOne (thesmartone):

Seems like you messed up in the first step, but then put the correct value at the end .-.

OpenStudy (anonymous):

24+(-10)=14 -4+4=0 (14,0)?

OpenStudy (anonymous):

yeah sorry that negative was an accident

OpenStudy (phi):

yes

TheSmartOne (thesmartone):

good job :)

OpenStudy (anonymous):

Thank you!